8 ,44 , 258 , 1540 ,...? What comes next and how?
Answer by Edwin McCravy(20056)

8 ,44 , 258 , 1540 ,...? What comes next and how?

                                 8  <-- begin with 8
   8×6 =   48, then   48- 4 =   44
  44×6 =  264, then  264- 6 =  258
 258×6 = 1548, then 1548- 8 = 1540
1540×6 = 9240, then 9240-10 = 9230

Edwin